1. Nile Blue in acid solution is a protonated imine with a pKa 10.1. Most protonated imines have pKa values of 7-8. raw structures for the conjugate acid of nile blue and use them to explain why it is less acidic than a typical imine.  
Answer #1

7 (Nile blue) U T =Reconance stabiliseel pka = 1001 R-CH=NA HO R-cn = NR Potonated Imone plus 7-8 I nile blue = is stasilised